A certain oscillator satisfies the equation of motion: 𝑥••+ 4x = 0. Initially the particle is at




the point x = √3 when it is projected towards the origin with speed 2.




2.1. Show that the position, x, of the particle at any given time, t, is given by:




x = √3 cos 2t – sin 2t.

1
Expert's answer
2022-03-02T14:38:22-0500

"\ddot x+4x=0,"

"x=c_1\cos 2t+c_2\sin 2t,"

"\dot x=2c_2\cos 2t-2c_1\sin 2t,"

"c_1=\sqrt 3,"

"c_2=1,"

"x=\sqrt 3\cos 2t+\sin 2t ."